OpenStudy (anonymous):

show that the sum of the elements of principal diagonal of (A+B)=sum of the elements of principal diagonal A+sum of the elements of principal diagonal B.....A and B are matrix thank you in advance

OpenStudy (jamesj):

They're both square matrices, yes?

OpenStudy (jamesj):

Write \( A = (a_{ij}) \) where A is a n x n matrix. The sum of the principal diagonal is \[ a_{11} + a_{22} + .... + a_{nn} \] and the some of the principal diagonal for B is a similar expression. Now, the sum of the two matrices \( A + B = (a_{ij} + b_{ij}) \) So it's diagonal elements are just \( a_{11} + b_{11}, a_{22} + b{22}, ... \)

OpenStudy (jamesj):

So it should now be obvious why this the relation asked is true.

OpenStudy (jamesj):

If you're having a hard time with this, write out generic 3x3 matrices for A and B and see how it works.

OpenStudy (anonymous):

yes they are both square matrices yes i know how it works but i don't know to express it in general temrs

OpenStudy (jamesj):

So technically the name for the sum of the principal diagonal is the trace. We write the trace of a n x n matrix A is \[ tr(A) = \sum_{i=1}^n a_{ii} \]

OpenStudy (jamesj):

As the matrix \( A + B = (a_{ij} + b_{ij}) \), the trace \[ tr(A+B) = \sum_{i=1}^n (a_{ii} + b_{ii}) = \sum_{i=1}^n a_{ii} + \sum_{i=1}^n b_{ii} = tr(A) + tr(B) \]

OpenStudy (anonymous):

so its the same thing with these tr(kA)=ktrA tr(A^T(transposed))=trA and tr(A*B)=tr(B*A)

OpenStudy (jamesj):

Yes. Just be careful with the last one.

OpenStudy (anonymous):

prove that if A and B are antisymetric matrix than these are antisymetric too A^T ,A+B,A-B and k*A .please